Show that if B>0, then the graph of x²+Bxy=F, is a hyperbola if F≠0, and two intersecting lines if F=0.

1
Expert's answer
2021-12-09T13:52:23-0500

A conic in the two-dimensional plane is the solution set of an equation of the form


"ax^2 + bxy + cy^2 + dx + ey +f = 0"

for real parameters "a,b,c,d,e" and "f", where at least one of "a,b" and "c" is not zero.

If "b^2-4ac>0," then the graph is a hyperbola (or two intersecting lines).

We have


"x^2+Bxy=F"

"a=1, b=B, c=0, f=-F"

"b^2-4ac=B^2-4(1)(0)=B^2>0, B\not=0"

Therefore the graph of "x^2+Bxy=F," is a hyperbola or two intersecting lines., if "B\not=0."

If "F\not=0," we have a hyperbola


"y=-\dfrac{x}{B}+\dfrac{F/B}{x}"

If "F=0,x^2+Bxy=0," we have two intersecting lines

"x=0\ or\ y=-\dfrac{1}{B}x"